the remainder obtained when the expression 1!+2!+3!+.........+(95)! is divided by 15 is???????


PLS EXPLAIN!!!!

Grade:9

2 Answers

Ashwin Muralidharan IIT Madras
11 years ago

Hi Shivam,

 

Starting from 5! and furhter all the terms are divisible by 15.

So the given expression is 1! + 2! + 3! + 4! + 15K = 1+2+6+24+15K = 33+15K.

Hence the remainder will be 3, when the above is divided by 15.

 

Hope it helps.

 

Wish you all the best.

Regards,

Ashwin (IIT Madras).

Aditya Kartikeya
11 years ago

sir , 

its quite an interesting que but i dint  exactly get what u said!!

could u explain again!